# Fusion of JWST data – Demonstrating practical feasibility

### 05 Mar 2026 v1

@@ Appendix A Data preprocessing
The SyFu algorithm includes three pre-fusion steps to compute the forward operators, to co-register the pairs of input data and to cross-calibrate the operators. These stages are detailed in what follows. To lighten the notations, the three-dimensional data cube have been unfold into matrices whose rows (resp. columns) are associated with the spectral (resp. spatial) dimension.
@@
The data sets for d203-506 and Titan are extracted from the MAST archive, hence they already underwent a complete, up to stage 3, JWST science calibration pipeline processing (JWSTpipeline). This pipeline converts raw observations acquired with an up-the-ramp strategy into science-ready data with common astrophysical units (e.g., MJy/sr). It computes the coordinates, corrects bad pixels and cuts off most of the instrumental noise and cosmic rays present in the data.
@@
where $L_{\mathrm{h}}$ and $W_{\mathrm{h}}(\cdot)$ denote the NIRSpec throughputs and convolutions by the PSF. The matrix $S$ is a regular subsampling matrix that models the difference in resolutions between the NIRCam and NIRSpec data.
@@
The NIRSpec cubes are resampled in the JWST pipeline so that the pixels are aligned with North. Since the NIRSpec aperture is not aligned with North at the moment of observation, this implies that the FoV is rotated. To obtain a rectangular aperture aligned with North we rotate the NIRSpec cube with the appropriate angle. We then apply a rotation to the NIRCam image, so that it has the same rotation with respect to the North as the NIRSpec cube. The next step consists in aligning the NIRCam images and NIRSpec cube. To do so, two cases need to be considered: non-moving (distant) objects and solar system objects. In the first case of non-moving objects, applicable to d203-506, we rely on the coordinates to align the NIRCam images and the NIRSpec cube. Finally, the NIRCam images are resampled to a pixel size that is 1/3 of NIRSpec (0.033”). This division by a factor of 3 is chosen since it is the closest integer to the pixel scale ratio between NIRSpec and NIRCam SW (3.2). Having an integer value provides significant simplifications and speeding-up of the algorithmic implementation (wei2015fast). In the second case of moving targets, applicable to Titan, the translations to be operated for the image alignment cannot be deduced from the pointing information. The NIRCam images and NIRSpec cubes are resampled to a common pixel size of 1/3 of the NIRSpec pixel size (0.033”). Then, using the phase cross-correlation technique (guizar2008crosscorrelation; scikit-image) we compensate for the translations. We note that all mentioned rotation and re-sampling steps were conducted with bicubic interpolations.
@@
Appendix B Fusion as a regularised inverse problem
@@
where $Z$ is the representation of the solution in the subspace and ${\Upsilon_{\mathrm{h}}}=\left[\overline{Y_{\mathrm{h}}},\ldots,\overline{Y_{\mathrm{h}}}\right]$ is a matrix whose columns are the mean spectrum $\overline{Y_{\mathrm{h}}}$ obtained by spatially averaging the NIRSpec input data. Beyond promoting spectrally consistent solutions, another key advantage of adopting the hardcore regularization (8) is that it allows the initial regularised least squares problem (3) to be reformulated with respect to the representation $Z$ of the solution in the subspace, significantly decreasing the computational complexity of the optimisation procedure.
@@
where $D$ is the matrix standing for the first-order finite difference operator and $\|\cdot\|_{\mathrm{F}}^{2}$ denotes the squared Frobenius norm, that is the sum of the squared coefficients of the matrix, the parameter $\mu$ adjusts the weight of the regularization. Since the columns of $V$ are orthogonal, this choice amounts to defining the spatial regularisation directly on the fused hyperspectral cube $X$ (guilloteau2020hyperspectral).
@@
It is worth noting that, thanks to the cross-calibration and the unit normalization of the PSFs previously discussed, the identity (6) ensures that $Y_{\mathrm{m}}-L_{\mathrm{m}}W_{\mathrm{m}}(\Upsilon_{\mathrm{h}})$ have a near zero mean.
@@
Computing the PSFs with WebbPSF takes several hours. They are computed once and then stored to avoid unnecessary re-computation. This can be done for each of the four NIRSpec filter/disperser combinations. To further lighten the computational burden, the associated large convolutional operators $W_{\mathrm{m}}(\cdot)$ and $W_{\mathrm{h}}(\cdot)$ are converted into term-wise multiplications in the Fourier domain. To do so, the associated PSF, as well as the NIRCam and NIRSpec data, are symmetrically padded and evaluated in the Fourier domain using a two-dimensional fast Fourier transform. This formulation enforces periodic boundary conditions, which may result into unwanted boundary effects, e.g., when bright sources are present near the edges of the FoV. Denoting by $h$ and $w$ the spatial height and width of the NIRSpec cube $Y_{\mathrm{h}}$, the size of the padded NIRSpec cube and PSF is $3h\times 3w$, while it is $9h\times 9w$ for NIRCam images and PSF. This choice preserves the cross-calibration, ensuring that both the original and padded data and PSF have the same mean.
Then, all quantities defined in (9) are vectorised adopting a sparse matrix representation. By setting the gradient of the resulting objective function to zero, the linear system to be solved is characterised by a large yet structured and highly sparse matrix that can be conveniently represented using the compressed sparse row format (CSR) of the Python SciPy library. Thanks to the highly sparse structure of the linear problem, this vectorization yields a computational complexity of $\mathcal{O}({p_{\mathrm{m}}k^{2}})$ for each iteration of the subsequent iterative algorithm, where $p_{\mathrm{m}}$ stands for the number of NIRCam pixels (guilloteau2020hyperspectral). The matrices associated with this sparse linear system, whose expressions are detailed in (guilloteau2020hyperspectral), are computed once and stored for subsequent use. This computation takes less than one minute on a laptop when using a reasonably small subspace (here $k=3$, see Table 1).
Finally, the linear system is solved using the conjugate gradient descent implemented in the scipy.sparse.linalg library, which eases and fastens the tuning of the parameters $\mu$ and $\gamma$ (see Table 1). By initializing the algorithm with an interpolated counterpart of the NIRSpec data $Y_{\mathrm{h}}$, 1000 iterations of the algorithm are required to obtain satisfactory results (see also the code publicly available, as detailed in Appendix E).
@@
Manual tuning – The parameter $k$, which corresponds to the number of principal components to be kept when conducing the PCA, is adjusted to preserve sufficient information in the NIRSpec data. Beyond its impact on the spectral regularisation, it is worth noting that this parameter has also an impact on the dimension of the linear system to be solved. Thus, choosing a small $k$ reduces drastically the computational time required in the iterative minimization procedure. Regarding the parameter $\gamma$, it has been manually adjusted to reach a trade-off between the two data fitting terms. Finally, the weight of the regularisation $\mu$ is adjusted by performing a grid search with a few values between 1 and $10^{-4}$ and choosing the best one through visual inspection. Table 1 summarises the values of these parameters manually adjusted for the two data sets.
Automatic tuning – Techniques commonly adopted in image processing can be devised to automatically guide the selection of the hyperparameters $k$, $\gamma$ and $\mu$. One strategy consists in first estimating the noise levels ${\sigma}_{\mathrm{m}}$ and ${\sigma}_{\mathrm{h}}$ in the NIRCam and NIRSpec data, e.g., using the robust estimator proposed by donoho1995noising. Then, when implementing the hard spectral regularization, only the principal components with corresponding eigenvalues higher than $\sigma_{\mathrm{h}}^{2}$ should be kept in $V$, explicitly adjusting the number $k$. Finally, a straightforward interpretation of the two other hyperparameters can be offered by adopting an empirical Bayesian formulation of the regularised least-square problem (9). This interpretation leads to $\gamma={\sigma_{\mathrm{m}}^{2}}/{\sigma_{\mathrm{h}}^{2}}$ and $\mu={\sigma_{\mathrm{m}}^{2}}/{\alpha^{2}}$ where $\alpha^{2}$ is the mean of the squared Frobenius norm of $ZD$, that can be empirically estimated on a crude solution $\hat{Z}$, e.g., computed from an interpolated counterpart of the NIRSpec data. This strategy relies on the Gaussian assumption of the residual noises, which is not realistic because of the complex nature of the JWST NIRCam and NIRSpec data. Yet, this simple and fast strategy yields results similar to those reported in the paper.
Appendix C Data properties and range of symmetric fusion
@@
Appendix D Consistency of the fused hyperspectral cubes
@@
Figure 7: Comparison between the JWST NIRSpec and the fused data cube averaged spectra. a, d203-506 protoplanetary disk averaged NIRSpec spectrum (orange) and fused hyperspectral cube averaged spectra (dark blue). Relative error (in percent) between them is shown below, in blue. b same as a for Titan.
D.1 Measures of quality
Figure 8: Comparison between the JWST NIRCam and NIRSpec Titan data and the hyperspectral cube resulting from their fusion. a, images of Titan observed by NIRCam filters F182M, F200W and F210M (top) and the Titan fused hyperspectral cube integrated over the corresponding throughputs (bottom). b, NIRSpec averaged spectrum (orange), the fused hyperspectral cube average spectrum (dark blue). Relative error (in percent) between them is shown below, in blue.
@@ D.1 Measures of quality
The fused hyperspectral cubes (Fig. 4) are spectrally degraded using the NIRCam forward model in equation (4), producing $F$ images denoted by $X_{\mathrm{m},f}$ for $f\in[\![1;F]\!]$. We then compare these reconstructed images with the NIRCam measurements $Y_{\mathrm{m},f}$, by means of two metrics widely used in image processing. The first one, denoted by peak signal-to-noise ratio (PSNR), computes a normalised relative error between images $X$ and $Y$:
$\mathrm{PSNR}(Y,X)=10\mathrm{log}_{\rm 10}\frac{\max{(Y)}^{2}}{\mathrm{MSE}(Y,X)},$ (10)
where $\mathrm{MSE}(\cdot,\cdot)$ denotes the mean squared error. The better the reconstruction, the higher the $\mathrm{PSNR}$. In addition, we compute the structural similarity index (wang2004image), which is a local measure of similarity between two image patches $X$ and $Y$, designed to match human perception:
@@
where $\sigma(\cdot,\cdot)$ is the covariance, $\sigma(\cdot)$ is the standard deviation, $c_{1}=(0.01(\max(Y)-\min(Y))^{2}$ and $c_{2}=(0.03(\max(Y)-\min(Y))^{2}$ are two small variables stabilizing the division (wang2004image). The SSIM is computed locally on patches of size $7\times 7$, and then averaged over all patches in the image. The better the reconstruction, the higher the $\mathrm{SSIM}$ with a maximum of $1$. The numerical values are reported in Table 4.
@@
Data set Filter PSNR (dB) SSIM
@@
D.2 Consistency with the input
The consistency of the SyFu method can be evaluated by conducting a sanity check which assesses the spatial and spectral consistency of marginal products derived from the fused hyperspectral cubes with respect to the input NIRCam and NIRSpec data. More precisely, we first integrate the fused hyperspectral cubes over the throughputs of the NIRCam filters used for fusion. These images extracted from the fused cube are compared to the corresponding input NIRCam images in Fig. 6. The extracted images match very well the NIRCam images. This is also supported by the results on Table 4, where all PSNR values are $\gtrsim$ 30 dB (resp $\gtrsim$ 40 dB) and all SSIM values are >> 0.9 (resp. $\gtrsim$ 0.99) for d203-506 data (resp. Titan data), which is recognised as a good pixel-wise and perceptual similarity in the image processing literature.
@@ D.2 Consistency with the input
D.3 Additional validation tests
To further assess the consistency of the data fusion method, we considered the previously described Titan data set and rerun the fusion, but excluding NIRCam data for filters F187N and F212N. Figure 8a, b shows the marginal products derived from the resulting fused hyperspectral cube (as in Fig. 6 and 7) which demonstrate it pass the sanity check described in the Results section. We further validate the consistency of the fused hyperspectral cube in Fig. 9 where the unused F187N and F212N images are compared with the fused hyperspectral cube integrated over their respective throughputs. From a visual inspection, those images match each others closely, which is supported by respective PSNR values of 34.03 and 28.43 (see definition below). To conclude, the SyFu algorithm can generate hyperspectral cubes that are fully consistent even with NIRCam observations excluded from the fusion.
Appendix E Data availability

### 27 Mar 2026 v2

@@ Appendix B Data pre-processing
The SyFu algorithm includes three pre-fusion steps to compute the forward operators, to co-register the pairs of input data and to cross-calibrate the operators. These stages are detailed in the following sections. To lighten the notations, the three-dimensional data cube was unfolded into matrices whose rows (resp. columns) are associated with the spectral (resp. spatial) dimension.
@@
The datasets for d203-506 and Titan are extracted from the MAST archive, hence they already underwent a complete, up to stage 3, JWST science calibration pipeline processing (JWSTpipeline). This pipeline converts raw observations acquired with an up-the-ramp strategy into science-ready data with common astrophysical units (e.g. MJy/sr). It computes the coordinates, corrects bad pixels and cuts off most of the instrumental noise and cosmic rays present in the data.
@@
where $L_{\mathrm{h}}$ and $W_{\mathrm{h}}(\cdot)$ denote the NIRSpec throughputs and convolutions by the PSF. The matrix $S$ is a regular sub-sampling matrix that models the difference in resolutions between the NIRCam and NIRSpec data.
@@
The NIRSpec cubes are resampled in the JWST pipeline so that the pixels are aligned with North. Since the NIRSpec aperture is not aligned with North at the moment of observation, this implies that the FoV is rotated. To obtain a rectangular aperture aligned with North we rotate the NIRSpec cube with the appropriate angle. We then apply a rotation to the NIRCam image, so that it has the same rotation with respect to the North as the NIRSpec cube. The next step consists in aligning the NIRCam images and NIRSpec cube. To do so, two cases need to be considered: non-moving (distant) objects and solar system objects. In the first case of non-moving objects, applicable to d203-506, we rely on the coordinates to align the NIRCam images and the NIRSpec cube. Finally, the NIRCam images are resampled to a pixel size that is one-third of NIRSpec (0.033”). This division by a factor of three is chosen since it is the closest integer to the pixel scale ratio between NIRSpec and NIRCam short wavelength channel (3.2). Having an integer value provides significant simplifications and speeding-up of the algorithmic implementation (wei2015fast). In the second case of moving targets, applicable to Titan, the translations to be operated for the image alignment cannot be deduced from the pointing information. The NIRCam images and NIRSpec cubes are resampled to a common pixel size of one-third of the NIRSpec pixel size (0.033”). Then, using the phase cross-correlation technique (guizar2008crosscorrelation; scikit-image) we compensate for the translations. We note that all mentioned rotation and re-sampling steps were conducted with bicubic interpolations.
@@
Appendix C Fusion as a regularised inverse problem
@@
where $Z$ is the representation of the solution in the subspace and ${\Upsilon_{\mathrm{h}}}=\left[\overline{Y_{\mathrm{h}}},\ldots,\overline{Y_{\mathrm{h}}}\right]$ is a matrix whose columns are the mean spectrum $\overline{Y_{\mathrm{h}}}$ obtained by spatially averaging the NIRSpec input data. Beyond promoting spectrally consistent solutions, another key advantage of adopting the hardcore regularisation (8) is that it allows the initial regularised least squares problem (3) to be reformulated with respect to the representation $Z$ of the solution in the subspace, significantly decreasing the computational complexity of the optimisation procedure.
@@
where $D$ is the matrix standing for the first-order finite difference operator and $\|\cdot\|_{\mathrm{F}}^{2}$ denotes the squared Frobenius norm, that is the sum of the squared coefficients of the matrix, the parameter $\mu$ adjusts the weight of the regularisation. Since the columns of $V$ are orthogonal, this choice amounts to defining the spatial regularisation directly on the fused hyperspectral cube $X$ (guilloteau2020hyperspectral).
@@
It is worth noting that, thanks to the cross-calibration and the unit normalisation of the PSFs previously discussed, the identity (6) ensures that $Y_{\mathrm{m}}-L_{\mathrm{m}}W_{\mathrm{m}}(\Upsilon_{\mathrm{h}})$ have a near zero mean.
@@
Computing the PSFs with WebbPSF takes several hours. They are computed once and then stored to avoid unnecessary re-computation. This can be done for each of the four NIRSpec filter and disperser combinations. To further lighten the computational burden, the associated large convolutional operators $W_{\mathrm{m}}(\cdot)$ and $W_{\mathrm{h}}(\cdot)$ are converted into term-wise multiplications in the Fourier domain. To do so, the associated PSF, as well as the NIRCam and NIRSpec data, are symmetrically padded and evaluated in the Fourier domain using a two-dimensional fast Fourier transform. This formulation enforces periodic boundary conditions, which may result into unwanted boundary effects, for example, when bright sources are present near the edges of the FoV. Denoting by $h$ and $w$ the spatial height and width of the NIRSpec cube $Y_{\mathrm{h}}$, the size of the padded NIRSpec cube and PSF is $3h\times 3w$, while it is $9h\times 9w$ for NIRCam images and PSF. This choice preserves the cross-calibration, ensuring that both the original and padded data and PSF have the same mean.
Then, all quantities defined in (9) are vectorised adopting a sparse matrix representation. By setting the gradient of the resulting objective function to zero, the linear system to be solved is characterised by a large yet structured and highly sparse matrix that can be conveniently represented using the compressed sparse row format (CSR) of the Python SciPy library. Thanks to the highly sparse structure of the linear problem, this vectorisation yields a computational complexity of $\mathcal{O}({p_{\mathrm{m}}k^{2}})$ for each iteration of the subsequent iterative algorithm, where $p_{\mathrm{m}}$ stands for the number of NIRCam pixels (guilloteau2020hyperspectral). The matrices associated with this sparse linear system, whose expressions are detailed in (guilloteau2020hyperspectral), are computed once and stored for subsequent use. This computation takes less than one minute on a laptop when using a reasonably small subspace (here $k=3$, see Table 1).
Finally, the linear system is solved using the conjugate gradient descent implemented in the scipy.sparse.linalg library, which eases and fastens the tuning of the parameters $\mu$ and $\gamma$ (see Table 1). By initialising the algorithm with an interpolated counterpart of the NIRSpec data $Y_{\mathrm{h}}$, 1000 iterations of the algorithm are required to obtain satisfactory results (see also the code publicly available, as detailed in Appendix A).
@@
Manual tuning – The parameter $k$, which corresponds to the number of principal components to be kept when conducing the PCA, is adjusted to preserve sufficient information in the NIRSpec data. Beyond its impact on the spectral regularisation, it is worth noting that this parameter has also an impact on the dimension of the linear system to be solved. Thus, choosing a small $k$ reduces drastically the computational time required in the iterative minimisation procedure. Regarding the parameter $\gamma$, it has been manually adjusted to reach a trade-off between the two data fitting terms. Finally, the weight of the regularisation $\mu$ is adjusted by performing a grid search with a few values between 1 and $10^{-4}$ and choosing the best one through visual inspection. Table 1 summarises the values of these parameters manually adjusted for the two datasets.
Automatic tuning – Techniques commonly adopted in image processing can be devised to automatically guide the selection of the hyperparameters $k$, $\gamma$ and $\mu$. One strategy consists in first estimating the noise levels ${\sigma}_{\mathrm{m}}$ and ${\sigma}_{\mathrm{h}}$ in the NIRCam and NIRSpec data, for example, using the robust estimator proposed by donoho1995noising. Then, when implementing the hard spectral regularisation, only the principal components with corresponding eigenvalues higher than $\sigma_{\mathrm{h}}^{2}$ should be kept in $V$, explicitly adjusting the number $k$. Finally, a straightforward interpretation of the two other hyperparameters can be offered by adopting an empirical Bayesian formulation of the regularised least-square problem (9). This interpretation leads to $\gamma={\sigma_{\mathrm{m}}^{2}}/{\sigma_{\mathrm{h}}^{2}}$ and $\mu={\sigma_{\mathrm{m}}^{2}}/{\alpha^{2}}$ where $\alpha^{2}$ is the mean of the squared Frobenius norm of $ZD$, that can be empirically estimated on a crude solution $\hat{Z}$, for example, computed from an interpolated counterpart of the NIRSpec data. This strategy relies on the Gaussian assumption of the residual noises, which is not realistic because of the complex nature of the JWST NIRCam and NIRSpec data. Yet, this simple and fast strategy yields results similar to those reported in the paper.
Appendix D Data properties and range of symmetric fusion
@@
Appendix E Consistency of the fused hyperspectral cubes
@@
Figure 7: Comparison between the JWST NIRSpec and the fused data cube averaged spectra. a, d203-506 protoplanetary disk averaged NIRSpec spectrum (orange) and fused hyperspectral cube averaged spectra (dark blue). Relative error (in percent) between them is shown below, in blue. b : Same as a, but for Titan.
E.1 Measures of quality
Figure 8: Comparison between the JWST NIRCam and NIRSpec Titan data and the hyperspectral cube resulting from their fusion. a: Images of Titan observed by NIRCam filters F182M, F200W and F210M (top) and the Titan fused hyperspectral cube integrated over the corresponding throughputs (bottom). b, NIRSpec averaged spectrum (orange), the fused hyperspectral cube average spectrum (dark blue). Relative error (in percent) between them is shown below, in blue.
@@ E.1 Measures of quality
The fused hyperspectral cubes (Fig. 4) are spectrally degraded using the NIRCam forward model in equation (4), producing $F$ images denoted by $X_{\mathrm{m},f}$ for $f\in[\![1;F]\!]$. We then compare these reconstructed images with the NIRCam measurements $Y_{\mathrm{m},f}$, by means of two metrics widely used in image processing. The first one, denoted by peak signal-to-noise ratio (PS/N), computes a normalised relative error between images $X$ and $Y$:
$\mathrm{PS/N}(Y,X)=10\mathrm{log}_{\rm 10}\frac{\max{(Y)}^{2}}{\mathrm{MSE}(Y,X)},$ (10)
where $\mathrm{MSE}(\cdot,\cdot)$ denotes the mean squared error. The better the reconstruction, the higher the $\mathrm{PS/N}$. In addition, we compute the structural similarity index (wang2004image), which is a local measure of similarity between two image patches $X$ and $Y$, designed to match human perception:
@@
where $\sigma(\cdot,\cdot)$ is the covariance, $\sigma(\cdot)$ is the standard deviation, $c_{1}=(0.01(\max(Y)-\min(Y))^{2}$ and $c_{2}=(0.03(\max(Y)-\min(Y))^{2}$ are two small variables stabilising the division (wang2004image). The SSIM is computed locally on patches of size $7\times 7$, and then averaged over all patches in the image. The better the reconstruction, the higher the $\mathrm{SSIM}$ with a maximum of $1$. The numerical values are reported in Table 4.
@@
dataset Filter PS/N (dB) SSIM
@@
E.2 Consistency with the input
The consistency of the SyFu method can be evaluated by conducting a sanity check which assesses the spatial and spectral consistency of marginal products derived from the fused hyperspectral cubes with respect to the input NIRCam and NIRSpec data. More precisely, we first integrate the fused hyperspectral cubes over the throughputs of the NIRCam filters used for fusion. These images extracted from the fused cube are compared to the corresponding input NIRCam images in Fig. 6. The extracted images match very well the NIRCam images. This is also supported by the results on Table 4, where all PS/N values are $\gtrsim$ 30 dB (resp $\gtrsim$ 40 dB) and all SSIM values are >> 0.9 (resp. $\gtrsim$ 0.99) for d203-506 data (resp. Titan data), which is recognised as a good pixel-wise and perceptual similarity in the image processing literature.
@@ E.2 Consistency with the input
E.3 Additional validation tests
To further assess the consistency of the data fusion method, we considered the previously described Titan dataset and rerun the fusion, but excluding NIRCam data for filters F187N and F212N. Figure 8a, b shows the marginal products derived from the resulting fused hyperspectral cube (as in Fig. 6 and 7) which demonstrate it pass the sanity check described in the Results section. We further validate the consistency of the fused hyperspectral cube in Fig. 9 where the unused F187N and F212N images are compared with the fused hyperspectral cube integrated over their respective throughputs. From a visual inspection, those images match each other closely, which is supported by respective PS/N values of 34.03 and 28.43 (see definition below). To conclude, the SyFu algorithm can generate hyperspectral cubes that are fully consistent even with NIRCam observations excluded from the fusion.